CH/Tech/93/07

Amplifier Development

Schematic of Control Unit

Author: W.N. Catford Date: 4 April 1993

This diagram is my interpretation of the system described in Ch/Tech/93/05 by Dick Hunt. It contains no major errors, but any errors of detail are mine:

a